OpenStudy (kitten_is_back):

What is the value of the variable that makes the statement true? c + (–8) + (–5) = 0 A. –13 B. –3 C. 3 D. 13

OpenStudy (anonymous):

D

OpenStudy (kitten_is_back):

thanks :p

OpenStudy (anonymous):

13-8=5 5-5=0

OpenStudy (anonymous):

try plugging each one in, -13 + (-8) + (–5) = -26 so that can't be right -3 +(-8) + (–5) = -16 so wrong again 3 + (-8) + (–5) = -10 so no 13 + (-8) + (-5) = 0 so D is correct

OpenStudy (kitten_is_back):

oh oh i gt it thanks :D